1.the sum of a number and fourteen is less or equal to three times the number
2.negative three times a number increased by seven is less than negative eleven
3.five times a number decreased by eight is at most ten more than negative three
4.Seven more than five sixths of a number is more than negative three
5.four times the sum of a number and two increased by three is at least twenty seven

Answer by mananth(12270)


I will set up the equations for you. I think then you can work them out.
1.the sum of a number and fourteen is less or equal to three times the number
let number be x
x+14<=3x
2.negative three times a number increased by seven is less than negative eleven
let number be x
-3x+7 <-11
3.five times a number decreased by eight is at most ten more than negative three
5x-8<=10+-3
4.Seven more than five sixths of a number is more than negative three
5x/6 +7 >-3
5.four times the sum of a number and two increased by three is at least twenty seven
4(x+2)+3>=27
...
m.ananth@hotmail.ca